The perimeter of a square must be greater than 110 inches but less than 126 inches. Find the range of the possible side lengths that satisfy these conditions
Answer by Theo(13342)

p = 4s

p = perimeter of a square.
s equal side length of a square.

the perimeter must be greater than 110 but less than 126.

4s = 110 results in s = 110/4 = 27.5

4s = 126 results in s = 126/4 = 31.5

the length of each side of the square has to be greater than 27.5 but less than 31.5.